The base of a triangular sail is 15 meters shorter than its height.The area is 92m^2. Find the base and the height of the sail.
Answer by Cromlix(4381)

area of triangle = 1/2( base x height)
Height = xm
Base = (x - 15)m
Area = 1/2(x(x - 15))
92 = 1/2(x^2 - 15x)
Multiply both sides by 2
x^2 - 15x = 184
x^2 - 15x - 184 = 0
(x + 8)(x - 23) = 0
x + 8 = 0
x = -8 (no answer as negative)
x - 23 = 0
x = 23 (This is your answer)
1/2 (23(23-15))= 92m
1/2 (23(8)) = 92m
1/2(184) = 92m
Hope this helps.
:-)